OpenStudy (anonymous):

Find the expected number of heads. Three coins are tossed. Number of heads: 0, 1, 2, 3 Probability: 1/8, 3/8, 3/8, 1/8 Find the expected number of heads. I know how to do this but I am not good with fractions can you please solve it now? P= 0(1/8) +1(3/8) +2(3/8)+3(1/8)

OpenStudy (anonymous):

multiply. what is 0 x 1/8

OpenStudy (anonymous):

0

OpenStudy (anonymous):

1 x 3/8 = 2 x 3/8 = 3 x 1/8 =

OpenStudy (anonymous):

I came up with 1.5 for the answer but not sure if its correct.

OpenStudy (anonymous):

Correct. it also helps if you put it into decimals but if you have to answer the question in a fraction look at it as